### v2

# A Hybrid Mean Field Framework for Aggregators Participating in Wholesale Electricity Markets

Abstract
The rapid growth of distributed energy resources (DERs), including rooftop solar and energy storage, is transforming the grid edge, where distributed technologies and customer-side systems increasingly interact with the broader power grid. DER aggregators, entities that coordinate and optimize the actions of many small-scale DERs, play a key role in this transformation. This paper presents a hybrid Mean-Field Control (MFC) and Mean-Field Game (MFG) framework for integrating DER aggregators into wholesale electricity markets. Unlike traditional approaches that treat market prices as exogenous, our model captures the feedback between aggregators’ strategies and locational marginal prices (LMPs) of electricity. The MFC component optimizes DER operations within each aggregator, while the MFG models strategic interactions among multiple aggregators. To account for various uncertainties, we incorporate reinforcement learning (RL), which allows aggregators to learn optimal bidding strategies in dynamic market conditions. We prove the existence and uniqueness of a mean-field equilibrium and validate the framework through a case study of the Oahu Island power system. Results show that our approach reduces price volatility and improves market efficiency, offering a scalable and decentralized solution for DER integration in wholesale markets.
1   Introduction
The accelerating deployment of distributed energy resources (DERs), including rooftop solar ane energy storage, is altering the structure and dynamics of power systems. These technologies are typically located behind the meter; that is, on the customer’s side of the electricity meter, where energy is generated, stored, or managed locally rather than delivered by the utility. As a result, end-users are evolving from passive consumers into prosumers – entities that both produce and consume electricity. These prosumers are becoming increasingly active in grid operations through bi-directional power flows, demand flexibility, and localized generation, collectively transforming the traditionally centralized grid into a more distributed and participatory system.
To support this shift, the Federal Energy Regulatory Commission (FERC) issued Order $2$ in 2020, allowing DERs to participate in wholesale electricity markets. However, individual DERs are often too small to meet the size or performance thresholds set by independent system operators (ISOs). Aggregators, also known as virtual power plants (VPPs), now act as intermediaries. For example, OhmConnect aggregates over 200,000 households to bid into CAISO, Tesla has deployed aggregated Powerwall systems in ERCOT, and Swell Energy partners with Hawaiian Electric to operate an 80 MW residential battery VPP. Still, integrating large numbers of small, heterogeneous resources raises challenges for understanding how their decentralized participation affects market efficiency and system reliability.
Current research has extensively examined the strategies of single aggregators managing DER portfolios, often under the simplifying assumption that wholesale market prices, such as locational marginal prices (LMPs), are exogenous (for example, [1, 2]). While this perspective offers valuable insights into aggregator operations, it neglects an important aspect of market integration; that is, aggregators’ collective actions can influence LMPs. In practice, as aggregators participate in wholesale markets at scale, their decisions contribute to price formation, making the system inherently more coupled and dynamic. Modeling this price feedback is essential for capturing the full impact of DER integration on market outcomes.
To address this need, we propose a prescriptive, learning-based modeling framework grounded in mean-field game (MFG) theory. Rather than describing current market behavior, our approach aims to engineer a scalable and decentralized market structure in which aggregators learn to optimize their DER portfolios in response to market signals shaped by the collective behavior of others. Using reinforcement learning (RL), each aggregator adapts its strategy over time based on observed market feedback, enabling dynamic decision-making in uncertain and evolving conditions. We establish theoretical guarantees by proving the existence and uniqueness of a mean-field equilibrium (MFE), which characterizes the steady-state outcome when aggregators adopt policies learned through this process. By explicitly modeling the feedback between aggregator decisions and market prices, and enabling each aggregator to learn optimal bidding strategies in a decentralized and scalable way, our framework lays the groundwork for AI-enabled, automated control architectures that support real-time, market-integrated coordination of DERs, a defining characteristic of what we believe future smart energy systems must support.
While the MFG framework captures interactions among aggregators at the market level, each aggregator still faces the challenge of managing a large number of DERs internally. To ensure scalability and tractability, we employ a mean-field control (MFC) approach to optimize the operation of each aggregator’s DER portfolio. A key difficulty lies in the presence of energy storage, which links decisions across time, and various uncertainties, including renewable intermittency, load fluctuations, and market price variability, whose distributions are often unknown or difficult to model accurately. Traditional optimization-based methods struggle in this setting due to their reliance on full system observability, centralized coordination, or precise probabilistic modeling. In contrast, our approach combines MFC with RL to learn optimal control policies from interaction with the environment, without requiring prior knowledge of system dynamics or uncertainty distributions. By solving for an optimal policy for a representative DER that reflects the population’s aggregate behavior, this framework significantly reduces computational complexity and enables each aggregator to learn adaptive, forward-looking strategies in a decentralized and uncertain environment.
This paper makes three key contributions: (1) It develops a hybrid MFC–MFG framework, integrated with RL, to enable decentralized DER aggregators to participate in wholesale electricity markets under uncertainty and price feedback. (2) It proves the existence and uniqueness of an MFE under entropy-regularized RL, ensuring well-posedness and consistency in the infinite-agent limit. (3) It implements a two-phase RL algorithm for the hybrid MFC–MFE model and demonstrates its effectiveness in an Oahu Island case study, showing stabilized aggregator behavior and market prices, reduced price volatility, and lower system-wide costs – highlighting the promise of AI-enabled coordination at the grid edge.
The rest of the paper is organized as follows. Section II reviews the related literature, and Section III introduces the baseline electricity market model and describes how LMPs are determined. Section IV formulates the aggregators’ decision-making problem and defines an MFE rigorously. Existence and uniqueness of an MFE are also presented. Section V describes the details of a two-phase distributed learning algorithm. Section VI presents numerical results based on the Oahu Island power system to demonstrate convergence and system-level benefits. Finally, Section VII concludes the paper and outlines directions for future research.
2   Literature Review
The literature on bidding strategies in wholesale electricity markets spans extensive work on both single-agent and multi-agent scenarios, using either optimization-based and learning-based methodologies. For single-agent bidding, notable contributions include [3], which formulates a two-level optimization model, and [4], which applies deep reinforcement learning (RL) techniques. Similar approaches have been extended to the multi-agent context. For instance, [5] leverages linear programming to address multi-agent demand response bidding, while [6] introduces a multilayer stochastic agent-based model to examine wholesale market dynamics with renewable energy resources. Additionally, [7] proposes an auction framework that facilitates aggregator participation in wholesale markets without necessitating real-time intervention from the distribution system operator (DSO). Learning-based approaches, in contrast, are predominantly applied to peer-to-peer energy trading markets. For example, [8] employs multi-agent reinforcement learning (MARL) with multi-armed bandits in repeated auction scenarios for peer-to-peer energy transactions.
The outcomes of MARL in these market settings often correspond to Nash equilibrium solutions within game theory. However, as the number of agents increases, scalability issues arise, rendering the problem computationally intractable [9]. To address this limitation, the concept of mean-field equilibrium (MFE) has gained significant traction in recent research. By assuming an infinite population, the mean field method reduces the multi-agent problem to an interaction between a single representative agent and the overall distribution of the population’s average state – referred to as the mean field. This framework significantly enhances computational tractability and well approximates optimal agent behaviors, particularly in large populations.
Various studies have explored mean-field frameworks in different settings. For scenarios where system dynamics are known, [10] employs a dynamic programming approach in response to the mean field. Conversely, in cases where system dynamics are unknown, reinforcement learning has been applied. For example, [11] develops a general MFG framework using RL under competitive conditions, while [12] incorporates fictitious play within MFG to enable agents to learn during gameplay. While most literature assumes the availability of a mean-field oracle, [13] outlines a Sandbox learning algorithm for situations where mean-field information is not directly accessible to agents. Finally, [14] investigates the use of mean-field control (MFC) for approximating MARL problems involving cooperative, heterogeneous agents.
3   Wholesale Market Model and Locational Marginal Prices
To ensure the paper is self-contained and accessible to a broad audience, this section presents a standard wholesale electricity market model along with the formulation of LMPs. We also cite a key Lipschitz continuity property of the LMP function, established in the second author’s earlier work [10], which plays a central role in enabling the theoretical analysis developed in this paper.
Consider a wholesale electricity market operating over a transmission network with $N$ buses, $L$ transmission lines, and $G$ bulk generators. Each bus $n\in\mathcal{N}:=\{1,\ldots,N\}$ serves a group of $M_{n}$ agents, comprising $M_{n}^{p}$ prosumers and $M_{n}^{c}$ pure consumers without any local generation or storage capabilities. Each bulk generator $g\in\mathcal{G}:=\{1,\ldots,G\}$ is characterized by a cost function $C_{g}(\cdot)$ that captures the cost of electricity production. Let $G_{n}\subseteq\mathcal{G}$ denote the set of generators connected to bus $n$. We require the sets to be disjoint and collectively exhaustive; that is, $\cup_{n=1}^{N}G_{n}=\mathcal{G}$ and $G_{n}\cap G_{n^{\prime}}=\emptyset$ for any $n\neq n^{\prime}$.
Figure 1: Conceptual framework of a wholesale energy market with prosumer participation through virtual power plants with aggregators, shown as a static snapshot at a single timestep $t$.
The wholesale market operates over discrete time intervals, indexed by $t\in\{1,2,\ldots\}$, where each timestep corresponds to a fixed duration, such as an hour. An overview of the market operation is illustrated in Fig. 1. During each timestep, the system operator collects aggregate supply and demand bids and solves an optimization problem to determine the least-cost dispatch that satisfies system constraints. The marginal cost of supplying one additional unit of electricity at each node, exactly the LMPs, is derived from the dual variables associated with the power balance and transmission constraints.
On the demand side, let $d^{n}_{it}$ denote the net demand of prosumer $i$ at bus $n$, defined as total energy consumption minus own solar generation. If $d^{n}_{it}>0$, the prosumer is a net consumer; if $d^{n}_{it}<0$, the prosumer sells energy to the grid. This bidirectional interaction allows prosumers to participate flexibly in the energy market. Similarly, we use $d^{n}_{jt}\geq 0$ to denote the demand of consumer $j$ at bus $n$, who does not generate energy. The total demand at bus $n$ and time $t$ is:
$D^{n}_{t}=\sum_{i=1}^{M^{p}_{n}}d^{n}_{it}+\sum_{j=1}^{M^{c}_{n}}d^{n}_{jt}.$ (1)
To avoid technical complications from infeasible supply-demand imbalances, we assume total net demand is non-negative at each timestep; that is, $\sum_{n=1}^{N}D^{n}_{t}\geq 0,\quad\forall t=1,2,\ldots.$ This is reasonable under current DER penetration levels, where distributed generation does not yet exceed total system demand.
We now explain how the aggregator constructs real-time demand bids. Each bus $n$ is served by a single aggregator representing its connected prosumers, who do not participate directly in the wholesale market. At each timestep $t$, the aggregator updates a control policy using an MFC approach (detailed next) and distributes it to prosumers. Each prosumer then determines its action (e.g., charging or discharging) based on its state. The aggregator collects and aggregates these actions into a net demand bid submitted to the system operator. Once all bids are received, the operator clears the market and computes the LMPs. The decision sequence is shown in Fig. 2.
Figure 2: Temporal workflow of the energy market framework. Aggregators train and distribute policies to prosumers at the beginning of each timestep. Prosumers act based on these policies, and their schedules are aggregated and submitted as bids to the system operator, who clears the market and determines LMPs at the end of each timestep.
To explain how the LMPs are actually determined, we present below a simplified version of the economic dispatch (ED) problem that the system operator solves. While actual dispatch problems in practice are considerably more complex, this minimal formulation retains the essential constraints, including power balance, generation capacity, and transmission limits, and is sufficient to illustrate how prices are formed and how they relate to system demand. Importantly, our modeling framework does not depend on this simplification. One of its key strengths is that it requires no change to the centralized operations of ISOs. System operators may continue using their full-scale market-clearing processes as is. All learning and coordination are implemented at the aggregator level, making our framework fully compatible with existing market structures.
The simplified ED problem is given as follows:
$\min_{p_{1t},\cdots,p_{Gt}}\;\sum_{g=1}^{G}C_{g}(p_{gt})$ (2)
[…]
$\quad-\overline{F}_{l}\leq\sum_{n=1}^{N}\text{PTDF}_{ln}\left(\sum_{g\in G_{n}}p_{gt}-D^{n}_{t}\right)\leq\overline{F}_{l},\ \forall\;l\in\{1,\ldots,L\}$ $(\underline{\mu}_{lt},\overline{\mu}_{lt})$ (4)
$\quad 0\leq p_{gt}\leq\overline{p}_{g},\forall\;g\in\{1,\ldots,G\}$ $(\underline{\nu}_{gt},\overline{\nu}_{gt}).$ (5)
Here, the variables $p_{gt}$ represent the average power output of generator $g$ during time interval $t$. The parameter $\text{PTDF}_{ln}$ denotes the power transfer distribution factor for line $l$ and bus $n$, which reflects how power flows across the network; $\overline{F}_{l}$ represents the maximum allowable power flow on line $l$, and $\overline{p}_{g}$ represents the maximum generation capacity of generator $g$. In addition, the dual variables are listed on the right-hand-side of each constraint. The dual of the power balance constraint (3), $\lambda^{\text{HUB}}_{t}$, represents the so-called hub price, and $\underline{\mu}_{lt},\overline{\mu}_{lt},\underline{\nu}_{gt},\overline{\nu}_{gt}$ are the dual variables corresponding to line flow limits of each line $l$, and generator output limits of each generator $g$, respectively.
The LMP at bus $n\in\{1,\ldots,N\}$ and time $t$, denoted $\lambda^{n}_{t}$, is derived as follows:
$\lambda^{n}_{t}$ :-$\coloneq\frac{\partial\mathfrak{L}_{t}}{\partial D^{n}_{t}}=\lambda^{\text{HUB}}_{t}-\sum_{l=1}^{L}\text{PTDF}_{ln}(\underline{\mu}_{lt}-\overline{\mu}_{lt}),:$ (6)
where $\mathfrak{L}_{t}$ represents the Lagrangian function of the ED problem. For notation simplicity, let $\bm{D}_{t}:=(D^{1}_{t},\cdots,D^{N}_{t})$ denote the vector of all demand bids; then the LMP $\lambda^{n}_{t}$ is a function of $\bm{D}_{t}$, which we write as $\lambda^{n}_{t}(\bm{D}_{t})$.
In establishing the existence and uniqueness of an MFE in later sections, a key technical condition needed is the Lipschitz continuity of the mapping $\lambda^{n}_{t}(\bm{D}_{t})$. The following result, established in [10], provides sufficient conditions under which this property holds. For completeness, we restate it here, along with the required constraint qualification assumption.
Assumption 1.
(LICQ) Let $X(\bm{D}_{t})$ denote the feasible region of the economic dispatch problem defined in (2)–(5). We assume that, for all $t$ and all $\bm{D}_{t}$ such that $X(\bm{D}_{t})$ is non-empty, the linear independence constraint qualification (LICQ) holds at every feasible point in $X(\bm{D}_{t})$.
Proposition 1.
(Lipschitz Continuity of LMPs)[10] Given Assumption 1 and that each generator’s cost function $C_{g}(\cdot)$ is strongly convex and quadratic. Then, for all buses $n=1,\ldots,N$, the LMP $\lambda^{n}_{t}(\bm{D}_{t})$ is single-valued and Lipschitz continuous with respect to $\bm{D}_{t}$; that is, there exists a constant $L_{\lambda}>0$ such that for any $\bm{D}_{t}$ and $\widetilde{\bm{D}}_{t}\geq 0$, $|\lambda^{n}_{t}(\bm{D}_{t})-\lambda^{n}_{t}(\widetilde{\bm{D}}_{t})|\leq L_{\lambda}\|\bm{D}_{t}-\widetilde{\bm{D}}_{t}\|_{1}.$
4   Aggregators’ Problem and the Mean-Field Framework
In this section, we first formulate the individual decision-making problem faced by each aggregator. Acting on behalf of a large population of prosumers equipped with solar PVs and energy storage, each aggregator must determine control strategies for these distributed resources under multiple sources of uncertainty, including solar energy output, actual energy demand, and market prices. A key challenge lies in the presence of energy storage, which links decisions across time and leads to a high-dimensional stochastic dynamic optimization problem. To address this complexity in a scalable manner, especially since each aggregator typically manages a large number of prosumers, we adopt an MFC approach, which approximates the collective behavior of prosumers using a representative agent and enables decentralized learning of control policies.
Building on the MFC formulation at the aggregator level, we next define an MFE that captures the steady-state outcome of their dynamic interactions and establish conditions for its existence and uniqueness.
4.1   Components of An Aggregator’s Problem
We first introduce the key components of the game. Notation-wise, we use $x\sim\mathcal{Q}$ to indicate that $x$ follows the distribution $\mathcal{Q}$; we use $\mathcal{P}(\mathcal{X})$ to be the set of all Borel probability measures on any space $\mathcal{X}$; we use $|\mathcal{X}|$ to denote the cardinality of $\mathcal{X}$ if it is finite and discrete; and we use $[\bm{\mathrm{x}}]_{k}$ to denote the $k$-th entry of the vector $\bm{\mathrm{x}}$.
Time: In power systems, many processes follow strong diurnal patterns. For instance, electricity demand and solar generation vary significantly across hours of the day. Therefore, it is necessary to distinguish between different hours even if the state variables are otherwise identical. To capture this temporal structure, we define two mappings over the global timestep $t=0,1,2,\ldots$: let $H$ be the number of timesteps in a day, then
$T_{\text{hour}}(t)=t\bmod H\quad\text{and}\quad T_{\text{day}}(t)=\left\lfloor\frac{t}{H}\right\rfloor,$ (7)
which return the hour within the day and the day index, respectively. These mappings allow the model to encode intra-day temporal variation while treating each day as structurally similar.
Household agents (prosumers and consumers): With the $M^{p}_{n}$ prosumers at each bus $n$, each is assumed to be equipped with both solar PV and energy storage. The total aggregated storage capacity at each bus is assumed to be capped at $\overline{E}^{n}$. To capture heterogeneity among prosumers, we assume that the $M^{p}_{n}$ prosumers at each bus $n$ are divided into $K$ types or subgroups, indexed by $k\in\{1,2,\ldots,K\}$, based on their energy storage capacities, which are also associated with differing PV system sizes. Let $b^{n}_{k}\in[0,1]$ denote the proportion of type-$k$ prosumers at bus $n$, satisfying $\sum_{k=1}^{K}b^{n}_{k}=1$. To assign different average capacities to each type, we introduce a parameter $\theta^{n}_{k}>0$, which specifies the relative (unnormalized) capacity assigned to type-$k$ prosumers. For example, if $\theta_{2}=2\theta_{1}$, then type-2 prosumers have twice the storage capacity of type-1 prosumers. The resulting capacity of each type-$k$ prosumer $i$ at bus $n$ is then given as follows: $\overline{E}^{n}_{ik}:=\frac{\theta^{n}_{k}\overline{E}^{n}}{M^{p}_{n}\sum_{\kappa=1}^{K}\theta^{n}_{\kappa}b^{n}_{\kappa}}.$ The reason for defining each agent’s storage capacity this way reflects a key consideration in mean-field analysis, where the number of agents is infinite. To keep the total storage capacity at each bus finite in the limit, individual capacities must scale down accordingly. While real-world systems involve only finitely many agents, a mean-field model should serve as a good approximation when the population is large, since the impact of any single agent on aggregate outcomes becomes negligible.
Having defined individual storage capacities, we now express each prosumer’s net demand in normalized form. (For ease of notation, we omit the type-$k$ subindex for prosumers unless explicitly needed.) At time $t$, let $q^{n}_{it}$ denote the ratio of prosumer $i$’s net demand to its storage capacity $\overline{E}_{i}^{n}$, so that net demand is scaled to lie between $-1$ and $1$. This normalization facilitates the theoretical analysis without affecting generality. The ratio $q^{n}_{it}$ is a random variable, since net demand is inherently stochastic. We denote its distribution by $\mathcal{Q}^{p}_{n}$. However, neither the prosumers nor the aggregators are assumed to know this distribution. A key strength of our framework is that agents are not required to have any prior knowledge of the underlying distributions; they instead learn optimal behaviors through historical data and repeated interactions with the environment.
For the $M^{c}_{n}$ pure consumers at each bus $n$, to facilitate a unified analysis, we also assign each consumer a notional or a ‘reference storage capacity’ $\overline{E}_{j}^{n}$, and define $q^{n}_{jt}$ as the ratio of demand to this reference capacity. This ratio is also a random variable, whose distribution is denoted by $\mathcal{Q}^{c}_{n}$.
Aggregators: Although each prosumer could, in theory, learn an individual policy, this is not practical in reality, as most do not meet the 100 kW minimum threshold required by FERC Order No. $2$ to participate directly in the wholesale market. Instead, we adopt a mean-field approach, letting each aggregator at bus $n$ represent all prosumers at that location. While prosumers differ in PV and storage sizes, they share the same external conditions (e.g., prices, weather) at the same location; hence a common policy is justified. In our set up, we assume that prosumers’ solar generation is first used to meet local demand, with any surplus charged into storage automatically. The only decision involves how much additional energy to purchase for charging, or how much to discharge from storage and sell to the market. Each aggregator then learns a policy that maps the states (to be described in detail below) to storage actions (charging/discharging). The learned policy updated at each time $t$ (details provided in the next section), is then distributed to all prosumers, who execute actions accordingly. In this framework, aggregators serve as the learning agents, while prosumers simply follow the prescribed policy. Consumers, who lack energy storage, do not participate in the learning process.
Actions: As noted above, the only decisions in our setup involve charging/discharging energy storage. Let $\mathcal{A}\subseteq[-1,1]$ denote the action space for each aggregator, where each action $a^{n}_{t}\in\mathcal{A}$ represents the proportion of storage capacity to charge (if $a^{n}_{t}>0$) or discharge (if $a^{n}_{t}<0$) at time $t$, drawn according to the policy learned by the aggregator at bus $n$.
To ensure feasibility, we implement action masking, a common technique in RL that eliminates invalid actions based on the current state. Specifically, the algorithm assigns a value of $\infty$- ∞ to infeasible actions, resulting in zero probability of selection. For example, if the current storage level is 0.8 (that is, 80% full), any action $a>0.2$ or $a<-0.8$ would be masked out to prevent over-charging or over-discharging.
States: In our model, each aggregator at bus $n$ is characterized by three state variables at time $t$: (i) storage level $e^{n}_{t}\in[0,1]$ (as a percentage of capacity), (ii) net load $q^{n}_{t}$ (unrelated to storage charging/discharging), and (iii) the current hour of the day, $T_{\text{hour}}(t)$. Let the state space be $\mathcal{S}\subseteq\mathbb{R}^{3}$, where each element is a tuple of the form $s^{n}_{t}:=(e^{n}_{t},q^{n}_{t},T_{\text{hour}}(t))\in\mathcal{S}$.
Among the state variables, the net load $q^{n}_{t}$ is an exogenous random variable, mainly driven by solar irradiance and consumption behavior, and is unaffected by aggregators’ actions. Time of day is deterministic. The only state variable affected by decisions is the storage level, which evolves according to the following rule after action $a^{n}_{t}$:
$e^{n}_{t+1}:=\max\{\min\{e^{n}_{t}+a^{n}_{t},1\},0\}.$ (8)
Mean Field: With states and actions defined, the next step in standard RL (and Markov games) setups is to specify each agent’s reward function. In our setting, however, rewards do not depend on the individual actions of other agents but on the aggregate behavior of the population. Since we consider a large (infinite) number of agents, the influence of any single agent is negligible. Instead, each agent responds to the population-wide distribution of states and actions, known as the mean field. For simplicity and tractability, we begin by assuming that the state space $\mathcal{S}$ and action space $\mathcal{A}$ are discrete and finite. Let $\mathcal{L}^{n}_{t}\in\mathcal{P}(\mathcal{S}\times\mathcal{A})$ denote the mean field at bus $n$ and time $t$, where $\mathcal{P}(\mathcal{S}\times\mathcal{A})$ denotes the space of probability distributions over the joint state–action space $\mathcal{S}\times\mathcal{A}$. Its definition is given by
$\mathcal{L}^{n}_{t}(s,a)=\lim_{M^{p}_{n}\to\infty}\frac{1}{M^{p}_{n}}\sum_{i=1}^{M^{p}_{n}}\mathbbm{1}_{(s^{n}_{it},a^{n}_{it})=(s,a)},\quad\forall s\in\mathcal{S},\ a\in\mathcal{A},$ (9)
where $\mathbbm{1}_{(s^{n}_{it},a^{n}_{it})=(s,a)}=1$ if prosumer $i$ is in state $s$ and takes action $a$ at time $t$, and 0 otherwise. Intuitively, the mean field is the limit of the histogram of joint state–action pairs as the number of prosumers tends to infinity.
Reward: We now define the single-period reward function $r_{n}:\mathcal{S}\times\mathcal{A}\times\mathbb{R}^{H}\to\mathbb{R}$ for each aggregator $n$. Let $s^{n}_{t}$ and $a^{n}_{t}$ be the state and action at time $t$, and let $\bm{\lambda}^{n}:=(\lambda^{n}_{1},\dots,\lambda^{n}_{H})$ denote the LMP profile at bus $n$. The reward is given by
$r_{n}(s^{n}_{t},a^{n}_{t},\bm{\lambda}^{n})=-\lambda^{n}_{T_{\text{hour}}(t)}\cdot\overline{E}_{n}\cdot\big{(}\Phi(e^{n}_{t},a^{n}_{t},\eta_{n})+q^{n}_{t}\big{)},$ (10)
where $\eta_{n}\in(0,1]$ is the storage efficiency (assumed uniform across prosumers at bus $n$), and $\Phi(\cdot)$ adjusts the action for efficiency losses:
$\Phi(e,a,\eta)=\begin{cases}\max\{-e,a\}\cdot\eta,&\text{if }a<0\ \text{(discharging)},\\ \min\{1-e,a\}/\eta,&\text{if }a\geq 0\ \text{(charging)}.\end{cases}$ (11)
This adjustment accounts for efficiency losses during storage operations. For instance, if $a>0$ and $\min\{1-e,a\}=a$, then to increase the storage level by $a$ (as a percentage of total capacity), the aggregator must purchase $a/\eta_{n}$ worth of energy from the grid to achieve the desired charge level, reflecting the fact that storage inefficiency requires drawing more energy than is ultimately stored.
The reward represents net profit, computed as negative price multiplied by total energy demand. A positive total quantity, comprising net load $q^{n}_{t}$ and charging/discharging $\Phi$, indicates a net energy purchase (resulting in a negative reward), while a negative total quantity reflects energy sold back to the grid (resulting in a positive reward).
Policy and Entropy-Regularized Value Function: At each decision time $t$, aggregator $n$ learns a policy $\pi^{n}_{t}:\mathcal{S}\times\mathbb{R}^{H}\to\mathcal{P}(\mathcal{A}),$ which maps each state and LMP profile to a distribution over actions. Given the LMP profile $\bm{\lambda}^{n}\in\mathbb{R}^{H}$ and initial state $s^{n}_{0}$, the value of policy $\pi^{n}_{t}$ is defined as the expected sum of discounted rewards: $V_{n}(s^{n}_{0},\pi^{n}_{t},\bm{\lambda}^{n})=\mathbb{E}\left[\sum_{\tau=0}^{\infty}\gamma_{n}^{\tau}r_{n}(s^{n}_{\tau},a^{n}_{\tau},\bm{\lambda}^{n})\right],$ where $\gamma_{n}\in(0,1]$ is the discount factor, and the expectation is over the trajectory induced by the policy and external uncertainties.
To promote exploration and improve learning stability, we incorporate entropy regularization into the reward function. Specifically, the single-period reward is modified as:
$r^{\text{REG}}_{n}(s,a,\bm{\lambda}^{n})=r_{n}(s,a,\bm{\lambda}^{n})-\Omega(\pi(\cdot\mid s,\bm{\lambda}^{n})),$ (12)
where $\Omega(\cdot)$ is a $\rho$-strongly convex regularization function. A common choice is the negative entropy: $\Omega(\pi(\cdot\mid s,\bm{\lambda}^{n}))=\alpha\sum_{a\in\mathcal{A}}\pi(a\mid s,\bm{\lambda}^{n})\log\pi(a\mid s,\bm{\lambda}^{n}),$ with $\alpha>0$ controlling the strength of regularization. The corresponding regularized value function is:
$V^{\text{REG}}_{n}(s^{n}_{0},\pi^{n}_{t},\bm{\lambda}^{n})=\mathbb{E}\left[\sum_{\tau=0}^{\infty}\gamma_{n}^{\tau}r^{\text{REG}}_{n}(s^{n}_{\tau},a^{n}_{\tau},\bm{\lambda}^{n})\right].$ (13)
Each aggregator seeks to learn an optimal policy $\pi^{n*}_{t}$ that maximizes $V^{\text{REG}}_{n}$ at each time $t$.
4.2   Mean Field Equilibrium
An MFE is a fixed point in a dynamic system of infinitely many interacting agents, characterized by two key properties:(i) optimality, where each agent’s strategy is optimal given the mean field, and (ii) consistency, where the mean field coincides with the distribution of states and actions induced by all agents following that strategy.
Note that an MFE is not generally a Nash equilibrium, as the latter is typically defined for games with a finite number of agents. Whether an MFE arises as the limit of $M$-agent Nash equilibria as $M\to\infty$ is a nontrivial question. [15] provides sufficient conditions under which the infinite-agent MFE policy yields an $\epsilon$-Nash equilibrium for the corresponding finite-$M$ game. In this work, we focus on the properties of the MFE in the infinite-agent setting and leave its connection to finite-agent Nash equilibria to future research.
To define MFE formally, we introduce two operators: the optimality operator and the consistency operator. Let $\Pi:=\left\{\pi\mid\pi:\mathcal{S}\times\mathbb{R}^{H}\to\mathcal{P}(\mathcal{A})\right\}$ denote the space of state–dependent stochastic policies. The optimality operator is defined as follows.
Definition 1 (Optimality Operator).
For each aggregator $n$, the optimality operator $\Gamma^{n}_{1}:\mathcal{S}\times\mathbb{R}^{H}\to\Pi$ maps a state $s^{n}\in\mathcal{S}$ and LMP profile $\bm{\lambda}^{n}\in\mathbb{R}^{H}$ to an optimal policy $\pi^{n*}$; that is, $\pi^{n*}=\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}),$ such that for any policy $\pi^{n}\in\Pi$, $V^{\text{REG}}_{n}(s^{n},\pi^{n*},\bm{\lambda}^{n})\geq V^{\text{REG}}_{n}(s^{n},\pi^{n},\bm{\lambda}^{n}).$ Equivalently, the operator is defined as $\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}):=\arg\max_{\pi^{n}\in\Pi}V^{\text{REG}}_{n}(s^{n},\pi^{n},\bm{\lambda}^{n}).$
Note that the optimal policy $\pi^{n*}$ is not unique in general. However, our setup with strongly convex regularization term can ensure the uniqueness of optimal policy (see Proposition 2 in Appendix A.1). Next, we define both the consistency operator and the concept of consistency:
Definition 2 (Consistency Operator and Consistency).
For each aggregator $n$, the consistency operator is a mapping $\Gamma^{n}_{2}:\mathcal{P}(\mathcal{S}\times\mathcal{A})\times\Pi\to\mathcal{P}(\mathcal{S}\times\mathcal{A}),$ which takes an MF $\mathcal{L}^{n}\in\mathcal{P}(\mathcal{S}\times\mathcal{A})$ and a policy $\pi^{n}\in\Pi$, and returns the updated MF $\mathcal{L}^{n\prime}$ resulting from all prosumers at bus $n$ following policy $\pi^{n}$. A pair $(\mathcal{L}^{n},\pi^{n})$ satisfies the consistency condition if the mean field is a fixed point of the operator; that is, $\Gamma^{n}_{2}(\mathcal{L}^{n},\pi^{n})=\mathcal{L}^{n}.$
In our model, given the fixed LMP vector $\bm{\lambda}$, the consistency operator has the following form:
$\Gamma^{n}_{2}(\mathcal{L}^{n},\pi^{n})(s^{\prime},a^{\prime}):=\zeta\frac{1}{|\mathcal{S}||\mathcal{A}|}+(1-\zeta)\sum_{s,a}\mathcal{L}^{n}(s,a)\cdot P^{n}(s^{\prime}\mid s,a^{\prime})\cdot\pi^{n}(a^{\prime}\mid s,\bm{\lambda}),$ (14)
for all $s^{\prime}\in\mathcal{S},a^{\prime}\in\mathcal{A}$, and $\zeta\in(0,1)$ is the probability of a uniform noise into the MF update. $P^{n}(s^{\prime}\mid s,a^{\prime})$ denotes the state transition probability for the aggregator at bus $n$.
Let $\bm{\pi}=(\pi^{1},\cdots,\pi^{N})$ denote the policy profile of all aggregators. We now introduce the formal definition of an MFE.
Definition 3 (MFE).
A profile $\{(\pi^{n*},\bm{\lambda}^{n*})\}_{n=1}^{N}$ is a mean field equilibrium if, for each aggregator $n$ and any state $s^{n}\in\mathcal{S}$ at time $t$, the following conditions hold:
  1. 1.
    Optimality: Given $\bm{\lambda}^{n*}$, the optimal policy is obtained via the optimality operator: $\pi^{n*}=\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n*}).$
  2. 2.
[…]
The definitions of the optimality and consistency operators, along with that of an MFE, naturally lead to a three-step computational framework:
Step 1: Fix $\mathcal{L}^{n}_{t}$ and $\bm{\lambda}^{n}_{t}$. Given $s^{n}\in\mathcal{S}$, aggregator $n$ computes the optimal policy using the optimality operator: $\pi^{n*}_{t}=\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t})$.
[…]
Step 3: Repeat Steps 1 and 2 until $\bm{\lambda}^{n}_{t+1}$ converges to $\bm{\lambda}^{n}_{t}$.
The key idea behind the computational framework is that information about the mean field is embedded in the LMPs through the economic dispatch problem. At each timestep, aggregators update their policies based on observed LMPs, and the mean field evolves in response to these updated policies. While this defines a structured framework, it is not a complete algorithm: Step 1 assumes access to an optimal policy, yet both the LMPs in the reward function (10) and the transition kernel $P^{n}(s^{\prime}\mid s,a^{\prime})$, which depends on unknown net load distributions, are not known to the aggregators. Section 5 addresses this challenge by introducing an RL algorithm to learn these unknown quantities. For now, assuming Step 1 can be executed, we show below that the 3-step procedure forms a contraction mapping, leading to a unique fixed point, which is exactly an MFE.
4.3   Existence and Uniqueness of MFE under Regularization
Our approach here is constructive: we prove that that the proposed 3-step framework defines a contraction mapping over the space of policy and price profiles, ensuring convergence to a unique fixed point, which is exactly an MFE; thus establishing its existence and uniqueness. Before presenting the main result, we show that both the optimality and consistency operators are Lipschitz continuous. Proofs are provided in Appendix A.2 and A.3.
Theorem 1 (Lipschitz Continuity of $\Gamma^{n}_{1}$).
Under the same assumptions as in Proposition 1, in the regularized MFG, for any belief $\bm{\lambda}^{n},\bm{\lambda}^{n^{\prime}}$, there exists $L_{1}\geq 0$ such that:
$\sup_{s\in\mathcal{S}}\|\Gamma^{n}_{1}(s,\bm{\lambda}^{n})-\Gamma^{n}_{1}(s,\bm{\lambda}^{n^{\prime}})\|_{1}\leq L_{1}\|\bm{\lambda}^{n}-\bm{\lambda}^{n^{\prime}}\|_{1}.$ (15)
Theorem 2 (Lipschitz Continuity of $\Gamma^{n}_{2}$).
Under the same assumptions as in Proposition 1, for any MF $\mathcal{L}^{n},\mathcal{L}^{n^{\prime}}$ and policies $\pi^{n},\pi^{n^{\prime}}$, given the belief $\bm{\lambda}^{n}$, there exists $L_{2},L_{3}\geq 0$ such that
$\|\Gamma^{n}_{2}(\mathcal{L}^{n},\pi^{n})-\Gamma^{n}_{2}(\mathcal{L}^{n^{\prime}},\pi^{n})\|_{1}\leq L_{2}\|\mathcal{L}^{n}-\mathcal{L}^{n^{\prime}}\|_{1},$ (16)
$\|\Gamma^{n}_{2}(\mathcal{L}^{n},\pi^{n})-\Gamma^{n}_{2}(\mathcal{L}^{n},\pi^{n^{\prime}})\|_{1}\leq L_{3}\sup_{s\in\mathcal{S}}\|\pi^{n}(\cdot\mid s,\bm{\lambda}^{n})-\pi^{n^{\prime}}(\cdot\mid s,\bm{\lambda}^{n})\|_{1}.$ (17)
We now present the main theorem regarding the MFE existence and uniqueness.
Theorem 3 (Existence and Uniqueness of MFE).
Under the same LICQ and strongly convex quadratic cost function assumptions as in Proposition 1, given that $L_{1}L_{\text{MF}}L_{3}+L_{2}<1$, there exists a unique MFE following the 3-step procedure.
The specific Lipschitz constant $L_{\text{MF}}$ and the proof of the result are provided in Appendix A.4.
5   The Two-Phase RL Algorithm
We now turn to the key challenge of Step 1: computing the optimal policy $\pi^{n*}_{t}=\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t})$. To this end, we propose a two-phase, distributed mean-field RL algorithm executed within each time period $t$. Phase 1: Training – each aggregator independently trains its policy using an RL algorithm for a specified number of steps, based on either LMP forecasts or prior beliefs about the LMPs for time period $t$. Phase 2: Execution – all prosumers act based on their aggregator’s trained policy. Since the policies are stochastic, even identical prosumers may take different actions. These actions are then compiled by the aggregator into bids submitted to the market clearing process. LMPs for time period $t$ are then calculated by solving the ED problem. This two-phase process is illustrated earlier in Figure 2.This two-phase procedure is illustrated earlier in Figure 2.
Aggregators still do not have direct knowledge of external uncertainty distributions, such as those governing net load. However, a key feature of electricity markets can be leveraged: fluctuations in LMPs implicitly reflect both market dynamics and external randomness. Using a similar approach as in [10], we allow each aggregator to form and maintain its own belief about the LMPs for every timestep of the day, represented as a vector of length $H$. These beliefs, derived from historical observations, guide the aggregator’s RL-based decision-making in Phase 1. After the LMPs are realized through market clearing, aggregators update their beliefs adaptively.
We assume that all prosumers at the same bus share the LMP belief vector maintained by their aggregator, as the aggregator represents their collective interests and the actual LMP is identical for all agents at the same bus. Let $\hat{\bm{\lambda}}^{n}_{t}\in\mathbb{R}^{H}$ denote the belief vector maintained by aggregator $n$ at time $t$. Once the system operator solves the ED problem and returns the realized LMP $\lambda^{n}_{t}$ for bus $n$, the aggregator updates its belief according to:
$\hat{\bm{\lambda}}^{n}_{t+1}:=\hat{\bm{\lambda}}^{n}_{t}-\frac{\delta_{n}}{\sqrt{T_{\text{day}}(t)+1}}\left((\hat{\bm{\lambda}}^{n}_{t})^{\top}\mathbf{1}_{T_{\text{hour}}(t)}-\lambda^{n}_{t}\right)\mathbf{1}_{T_{\text{hour}}(t)}$ (18)
where $\delta_{n}\in[0.5,1]$ is a learning rate hyper-parameter for this LMP update rule, and $\mathbf{1}_{T_{\text{hour}}(t)}\in\mathbb{R}^{H}$ denotes the vector whose value is 1 at the $T_{\text{hour}}(t)$-th entry and 0 everywhere else. That is, the update rule changes only the $T_{\text{hour}}(t)$-th entry of the belief vector at each time $t$.
Training Phase: At the beginning of each time period $t$, each aggregator fixes its LMP belief and trains a policy using an RL algorithm, referred to generically as Alg (such as PPO, TRPO, or SAC), for $T_{\text{train}}$ steps.
Execution Phase: After each aggregator has learned a policy, this policy is distributed to its prosumers. At each bus $n$, each prosumer $i$ takes an action from the policy $a^{n}_{it}\sim\pi^{n*}_{t}(\cdot|s^{n}_{it},\hat{\lambda}^{n}_{t})$. Also, each prosumer $i$ and consumer $j$’s original net demand are realized following the distributions $\mathcal{Q}^{p}_{n}$ and $\mathcal{Q}^{c}_{n}$, respectively. The net demand in the quantity of energy now has the following forms for prosumer $i$ at bus $n$ at time $t$: $d^{n}_{it}=(\Phi(e^{n}_{t},a^{n}_{t},\eta_{n})+q^{n}_{t})\overline{E}^{n}_{i},$ and similarly, for each consumer $j$ at bus $n$ at time $t$: $d^{n}_{jt}=q^{n}_{jt}\overline{E}^{n}_{j}.$
Recall that we introduce a random regeneration probability $\zeta$, which gives each prosumer a positive probability of changing its storage capacity. This serves two purposes: it maintains a dynamic environment to encourage continual learning, even at a steady state, and it mimics real-world conditions where prosumers may enter or exit the system at a given location. The transition of the state of charge for each prosumer $i$ at bus $n$ from time $t$ to $t+1$ is defined as follows:
$e^{n}_{i,t+1}=\begin{cases}\text{Uniform}(0,1),\text{with probability $\zeta$},\\ \max\left\{\min\left\{e^{n}_{it}+a^{n}_{it},1\right\},0\right\},\text{with probability $1-\zeta$}.\end{cases}$ (19)
As a result, the aggregator at $n$ transition its storage level as follows:
$e^{n}_{t+1}=\frac{1}{\overline{E}_{n}}\sum_{i=1}^{M^{p}_{n}}e^{n}_{i,t+1}\overline{E}^{n}_{i},$ (20)
which is a weighted average of all its prosumers storage level. This storage level is then used as the initial storage level state to begin the RL training for time $t$ for $T_{\text{train}}$ steps. With this setup, we present the pseudo-code in Algorithm 1.
Input: Initial battery states $e^{n}_{0}\in[0,1]$, initial LMP beliefs $\hat{\bm{\lambda}}^{n}_{0}\in\mathbb{R}^{H}$ with learning rates $\delta_{n}\in[0.5,1]$, demand shapes $\mathcal{Q}^{p}_{n},\mathcal{Q}^{c}_{n}$; training step $T_{\text{train}}$; random regeneration probability $\zeta$; an RL algorithm Alg; time functions $T_{\text{hour}}(\cdot),T_{\text{day}}(\cdot)$.
for $t=0,1,\ldots$ do
       Training phase
       for Bus $n=1,\ldots,N$ do
             Train the aggregator for $T_{\text{train}}$ steps using Alg with initial storage $e^{n}_{t}$ under $\hat{\bm{\lambda}}^{n}_{t}$ to get $\pi^{n*}_{t}$ ;
       end for
       Execution phase
       Initialize an empty bid collector $D_{t}\leftarrow\{\}$ ;
       for Bus $n=1,\ldots,N$ do
             $D^{n}_{t}\leftarrow 0$ ;
             foreach Prosumer $i=1,\ldots,M^{p}_{n}$ do
                   Get net demand $q^{n}_{it}\sim\mathcal{Q}^{p}_{n}$ ;
                   Take actions $a^{n}_{it}\sim\pi^{n*}_{t}$ ;
                   Storage state transition with random regeneration as in (19) ;
             end foreach
            for Consumer $j=1,\ldots,M^{c}_{n}$ do
                   Get net demand $q^{n}_{jt}\sim\mathcal{Q}^{c}_{n}$ ;
             end for
            Compute next storage state $e^{n}_{t+1}$ as in (20) ;
       end for
      Solve $\text{ED}_{\lambda}$ and get $\lambda^{n}_{t}$ for all $n$, and update the LMP belief as in (18) ;
end for
Algorithm 1 A two-phase distributed mean-field RL algorithm with LMP beliefs and entropy regularization
6   Numerical Experiment
6.1   Test Case
We use the 37-bus synthetic network from [16], which reflects the geographical layout of Oahu, Hawaii, and includes the latitude and longitude of each bus. To better align this case with real-world conditions, we map each power plant operated by Hawaiian Electric, the sole utility provider on Oahu, to its nearest bus in the synthetic network, using data from [17]. The modified network includes 26 generators: 4 oil, 2 biomass, 17 utility-scale solar (distinct from household-level PVs), and 3 wind generators. We assume quadratic cost functions for the oil and biomass units, with coefficient values adapted from [18, 19] according to their respective generator types. The parameter ranges are summarized in Table 1, and each generator’s specific coefficients are uniformly sampled from these ranges.
Table 1: Non-renewable Generators Cost coefficients ($C(p)=ap^{2}+bp$)
Fuel Type $a\,(\$/\text{MW}^{2}\text{h})$ $b\,(\$/\text{MWh})$
Oil 0.0059 - 0.0342 19.98
Biomass 0.001 - 0.002 28.45 - 52.65
The generation cost for solar and wind units is set to zero. Their capacity factors are treated as random variables. To reflect temporal patterns without explicitly modeling time series dependencies, we use historical hourly capacity factor profiles as the expected values, with utility-scale solar data from [20] and wind from [21]. At each timestep, the average capacity factor is multiplied by an independent random noise factor: a triangular distribution $\Delta(0.8,1.2,1)$ for solar and $\Delta(0.5,1.5,1)$ for wind. Here, $\Delta(\alpha,\beta,m)$ denotes a triangular distribution with lower limit $\alpha$, upper limit $\beta$, and mode $m$. Figure 3 shows the resulting solar and wind capacity factor shapes.
Figure 3: Input shapes for solar and wind capacity factors. Shaded areas indicate the range of possible values due to random scaling.
Each bus in the network hosts two types of agents: 650 prosumers and 2,000 pure consumers. Prosumers are further categorized into three types, 500 small, 100 medium, and 50 large, equipped with storage capacities of 10, 20, and 30 kWh, respectively. For consistency, we assign a 10 kWh reference storage capacity to each pure consumer.
Daily demand profiles are based on hourly average data from [22]: prosumers use net demand (gross load minus DER generation), while consumers use gross load. The available dataset provides only a single standardized net load profile (as a percentage of capacity) for prosumers, which we apply to all three prosumer types. Consequently, larger storage capacities imply proportionally greater absolute solar generation and consumption.
To introduce demand uncertainty, we scale each agent’s average daily demand by a factor drawn from the triangular distribution $\Delta(0.8,1.2,1)$. Figure 4 shows the net demand profiles used for prosumers and consumers across the day.
Figure 4: Input net demand shape for prosumers and consumers as ratio to their energy storage capacity. Consumers are given a 10kWh reference storage capacity. Data is adapted from [22]. Shaded areas represent the noise bounds.
6.2   Results and Discussion
We employ PPO as the learning algorithm and set $T_{\text{train}}=1,200$, with $H=12$ (corresponding to 2-hour timesteps). The simulation is run for a period of 50 days and repeated 10 times using different random seeds. The experiments were conducted on a Windows 11 system equipped with a 13th Gen Intel(R) Core(TM) i7-13700KF (24 cores) and NVIDIA GeForce RTX 4070. Figure 5 show the hub prices over the first and last 3 days of the simulation. Each plot shows two curves: the orange dotted line represents the scenario without storage or RL, where prosumers follow the input net demand profile (that is, prosumers only have grid-tied solar PVs); the solid blue line shows the case with storage and RL, where prosumers adjust their charging/discharging according to the aggregators’ learned policies. At the start of the training and learning process, LMP hub prices in both scenarios are similar. As learning progresses, prices in the storage-and-RL scenario become less volatile than in the no-storage baseline. Additionally, the price curve with storage stabilizes and develops a consistent daily pattern, suggesting that a steady state has been reached.
Figure 5: Comparison of hub prices with and without energy storage over the first 3 days (top) and last 3 days (bottom) during training. Shaded areas show the one standard deviation error bounds across all simulations.
To better measure the impacts of storage and RL on price volatility, we adopt the incremental mean volatility (IMV) measure from [23] as the metric. The IMV of a sequence of LMPs $\{\lambda_{t}\}_{t=1}^{\infty}$ is defined as:
$IMV=\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}|\lambda_{t+1}-\lambda_{t}|.$ (21)
Figure 6 presents the IMV results over the last three days, comparing the two scenarios: with storage and RL, and without. The storage and RL scenario yields significantly lower IMVs with reduced price volatility.
Figure 6: Comparison of IMV of the last 3 days between two scenarios: with storage and RL, and without storage or RL. Shadow areas show the one standard deviation error bounds across all simulations.
We also compute the average daily cost for all prosumers and consumers, as shown in Figure 7. We observe that system-wide total daily costs are lower in the scenario with storage and RL for both prosumers and consumers. This demonstrates that our framework delivers tangible cost benefits to all users, not just prosumers, but also to traditional consumers.
Figure 7: Comparison of ex-post daily costs between two scenarios: with storage and RL, and without storage or RL, for prosumers (top) and consumers (bottom). Shaded areas show the one standard deviation error bounds across all simulations.
Figure 8 illustrates the system-wide average charging and discharging actions of prosumers over the final 3 days of simulation, averaged across 10 runs. The bar plots indicate prosumer actions at each timestep (positive for charging, negative for discharging), while the overlaid line shows the corresponding average storage level. A clear daily pattern emerges from the learning process: prosumers tend to charge their storage during late-night and midday hours and discharge during morning and evening hours. Both the action profile and storage level converge to a consistent daily cycle, indicating stabilization in agents’ behavior and state evolution.
Figure 8: System-average actions and storage level over the last 3 days across 10 simulations. Bars show average charging/discharging; the line shows average storage level. Prosumers tend to charge at night and midday, and discharge in the morning and evening. Both metrics converge to a stable daily cycle. Shaded areas and error bars indicate one standard deviation across simulations.
One well-known challenge in power systems with high solar penetration is the so-callled ‘duck curve’ – a net demand pattern characterized by steep ramping needs in the early evening when solar generation drops and consumption rises. This pattern strains grid flexibility and has become a major operational concern. Figure 4 shows that Oahu’s net load profiles for prosumers (without storage or coordinated control) exhibit a pronounced duck curve based on historical data
Our numerical results demonstrate that the proposed RL-based algorithmic framework effectively mitigates the duck curve issue. Figure 9 shows the average net demand over the final 10 days of simulation. The green shaded region represents the reshaped net demand under storage and learned charging/discharging strategies, while the overlaid original input profiles (the shaded blue area) correspond to the baseline scenario without storage or learning. The RL-based policy shifts energy consumption (charging) toward low-price periods, that is, midnight and midday, and reduces net demand during peak morning and evening hours, thereby flattening the duck curve and enhancing grid stability.
Figure 9: Comparison of final total net demand shapes: consumer and prosumer input shapes vs. prosumer demand shape with storage and RL. The vertical axis shows energy as a ratio to total storage capacity for each agent type. Shaded areas represent input noise bounds and one standard deviation over the last 10 days and all RL simulations.
7   Conclusion
In this paper, we engineer an algorithmic framework for market integration of DERs, where decentralized aggregators use automated RL algorithms to support the integration of DERs into wholesale electricity markets. Specifically, our approach enables aggregators to make intelligent charging and discharging decisions on behalf of their prosumers under various sources of uncertainty, including weather, demand, and market price fluctuations.
The core idea is to treat the LMPs as a mean-field signal, which encapsulates aggregate supply-demand dynamics and network constraints. Within the proposed hybrid MFC-MFG framework, we introduce a two-phase RL-based learning algorithm that allows each aggregator to iteratively learn optimal policies in a decentralized manner.
We provide solid theoretical foundations by establishing sufficient conditions under which an MFE exists and is unique, and show that our three-step solution approach converges to this equilibrium in an infinite-agent game. Numerical experiments demonstrate that incorporating storage with our RL-based framework significantly reduces LMP volatility, leading to a more stable market. Importantly, it also yields cost benefits for both prosumers and consumers, underscoring the broader value of market-drien DER integration.
Future work will explore richer bidding strategies, including demand response, improved EV modeling, and simulations at the scale of large power systems such as CAISO or ERCOT. Another important direction is to study the strategic behavior of aggregators or VPPs that recognize their influence on LMPs and may manipulate prices to their advantage. Such behavior could disadvantage other market participants and undermine long-term market stability, for example, by driving prices too low, discouraging future capacity investments, and ultimately raising concerns about system reliability.
Appendix A Proofs
A.1   Uniqueness of Optimal Policy under Regularization
Proposition 2.
The $\arg\max_{\pi}V^{\text{REG}}_{n}$ has a unique solution.
Proof.
The regularized value function $V^{\text{REG}}_{n}$, as defined in (13), consists of two terms: the expected reward $\mathbb{E}_{a\sim\pi^{n}}[r_{n}(s,a,\bm{\lambda})]$, and a strongly concave regularization term $-\Omega(\pi^{n})$. Since we assume a finite action set, the distribution $\pi^{n}$ corresponds to a probability mass function over a finite set, and the expectation is simply a weighted average of the rewards across actions. Therefore, the expected reward is linear in $\pi^{n}$. Then the sum of a linear function and a strongly concave function remains strongly concave. Hence, the maximizer $\arg\max_{\pi}V^{\text{REG}}_{n}$ is unique. ∎
A.2   Proof of Theorem 1
We first need properties of Fenchel conjugate from Lemma 15 in [24]:
Proposition 3.
[…]
Proof.
See Lemma 15 from [24]. ∎
We are now ready to present the proof to Theorem 1. Because the proof is symmetric for all aggregators, we drop the index $n$ for notation brevity.
Proof of Theorem 1.
We consider a family of reinforcement learning problems parameterized by the price vector $\bm{\lambda}$ (aka the LMPs), each corresponding to a distinct $Q$-function. Given $\bm{\lambda}$, we define
$Q_{\bm{\lambda}}(s,a)=r^{\text{REG}}(s,a,\bm{\lambda})+\gamma\sum_{s^{\prime}}Q_{\bm{\lambda}}^{\star}(s^{\prime})\,P(s^{\prime}\mid s,a),$ (22)
where $Q_{\bm{\lambda}}^{\star}(s):=\max_{a}Q_{\bm{\lambda}}(s,a)$ for each state $s\in\mathcal{S}$. Then, starting with the same policy $\pi$, for any $\bm{\lambda}_{1},\bm{\lambda}_{2}$, we have
$\|Q_{\bm{\lambda}_{1}}^{\star}-Q_{\bm{\lambda}_{2}}^{\star}\|_{\infty}$
$=\max_{s,a}\Big{|}r^{\text{REG}}(s,a,\bm{\lambda}_{1})-r^{\text{REG}}(s,a,\bm{\lambda}_{2})+\gamma\sum_{s^{\prime}}Q_{\bm{\lambda}_{1}}^{\star}(s^{\prime})P(s^{\prime}|s,a)-\gamma\sum_{s^{\prime}}Q_{\bm{\lambda}_{2}}^{\star}(s^{\prime})P(s^{\prime}|s,a)\Big{|}$
$\leq\max_{s,a}\Big{\{}|r(s,a,\bm{\lambda}_{1})-r(s,a,\bm{\lambda}_{2})|+\gamma\Big{|}\sum_{s^{\prime}}Q_{\bm{\lambda}_{1}}^{\star}(s^{\prime})P(s^{\prime}|s,a)-\sum_{s^{\prime}}Q_{\bm{\lambda}_{2}}^{\star}(s^{\prime})P(s^{\prime}|s,a)\Big{|}\Big{\}}$
$\leq L_{r}\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}+\gamma\|Q_{\bm{\lambda}_{1}}^{\star}-Q_{\bm{\lambda}_{2}}^{\star}\|_{\infty},$
where the first inequality holds by the triangular inequality and the fact that the regularized term $\Omega(\pi)$ cancels out when $\pi$ is the same. For the second inequality, the first term follows the Lipschitz continuity of the reward functions); the second term follows from the fact that $P(\cdot\mid s,a)$ is a probability distribution over $s^{\prime}$, so the absolute difference of weighted sums is bounded by the maximum difference of the weights. By rearranging the terms, we obtain the following result,
$\|Q_{\bm{\lambda}_{1}}^{\star}-Q_{\bm{\lambda}_{2}}^{\star}\|_{\infty}\leq\frac{L_{r}}{1-\gamma}\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}$ (23)
Note that by the definition of the regularized reward function in (12), Equation (22) can be rewritten as:
$Q_{\bm{\lambda}}(s,a)=r(s,a,\bm{\lambda})+\gamma\sum_{s^{\prime}}Q_{\bm{\lambda}}^{\star}(s^{\prime})P(s^{\prime}|s,a)-\Omega(\pi)=\langle q_{\bm{\lambda}}^{s},a\rangle-\Omega(\pi),$ (24)
where for any $s$,
$q_{\bm{\lambda}}^{s}:=r(s,\cdot,\bm{\lambda})+\gamma\sum_{s^{\prime}}Q_{\bm{\lambda}}^{\star}(s^{\prime})P(s^{\prime}|s,\cdot)$ (25)
which is $\left(L_{r}+\frac{\gamma L_{r}}{1-\gamma}\right)$-Lipschtiz continuous with respect to $\bm{\lambda}$, as shown below:
$\|q_{\bm{\lambda}_{1}}^{s}-q_{\bm{\lambda}_{2}}^{s}\|_{\infty}=\max_{a}\Big{|}r(s,a,\bm{\lambda}_{1})-r(s,a,\bm{\lambda}_{2})+\gamma\sum_{s^{\prime}}Q_{\bm{\lambda}_{1}}^{\star}(s^{\prime})P(s^{\prime}|s,a)-\gamma\sum_{s^{\prime}}Q_{\bm{\lambda}_{2}}^{\star}(s^{\prime})P(s^{\prime}|s,a)\Big{|}$
$\leq L_{r}\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}+\gamma\max_{a}\Big{|}\sum_{s^{\prime}}Q_{\bm{\lambda}_{1}}^{\star}(s^{\prime})P(s^{\prime}|s,a)-\sum_{s^{\prime}}Q_{\bm{\lambda}_{2}}^{\star}(s^{\prime})P(s^{\prime}|s,a)\Big{|}$
$\leq L_{r}\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}+\gamma\|Q_{\bm{\lambda}_{1}}^{\star}-Q_{\bm{\lambda}_{2}}^{\star}\|_{\infty}=\left(L_{r}+\frac{\gamma L_{r}}{1-\gamma}\right)\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1}.$ (26)
Finally, we apply the Fenchel conjugate property stated in Proposition 3. Recall that the optimal policy at state $s$, given belief $\bm{\lambda}$, is defined as $\Gamma_{1}^{n}(s,\bm{\lambda})=\nabla\Omega^{\star}(q_{\bm{\lambda}}^{s}),$ where $\Omega^{\star}$ is the Fenchel conjugate of the regularization function $\Omega$, and $q_{\bm{\lambda}}^{s}$ is the regularized soft Q-function defined in (25). Since $\Omega$ is $\rho$-strongly convex, its conjugate $\Omega^{\star}$ is $\frac{1}{\rho}$-smooth. This means its gradient is Lipschitz continuous with constant $\frac{1}{\rho}$, and thus for any $s\in\mathcal{S}$,
$\|\Gamma^{n}_{1}(s,\bm{\lambda}_{1})-\Gamma^{n}_{1}(s,\bm{\lambda}_{2})\|_{1}\leq\frac{1}{\rho}\|q_{\bm{\lambda}_{1}}^{s}-q_{\bm{\lambda}_{2}}^{s}\|_{\infty}\leq\frac{1}{\rho}\left(L_{r}+\frac{\gamma L_{r}}{1-\gamma}\right)\|\bm{\lambda}_{1}-\bm{\lambda}_{2}\|_{1},$ (27)
where the second inequality comes from (26). Then setting $L_{1}=\frac{1}{\rho}\left(L_{r}+\frac{\gamma L_{r}}{1-\gamma}\right)$ completes the proof. ∎
A.3   Proof of Theorem 2
We prove that the consistency operator $\Gamma^{n}_{2}$, as defined in (14), is Lipschitz continuous with respect to both the MF distribution $\mathcal{L}^{n}$ and the policy $\pi^{n}$.
Proof of Theorem 2.
The consistency operator in (14) consists of two components: (i) the random regeneration of an agent, which occurs with probability $\zeta$, and (ii) the state transition when regeneration does not occur. Since the regeneration step is identical across all agents and independent of the current distribution $\mathcal{L}^{n}$, we focus on the second component – the actual state transition. To simplify notation, we use $\tilde{\Gamma}_{2}^{n}$ to denote only the state transition dynamics:
$\tilde{\Gamma}_{2}^{n}(\mathcal{L}^{n},\pi^{n})(s^{\prime},a^{\prime})=\sum_{s,a}\mathcal{L}^{n}(s,a)\cdot P^{n}(s^{\prime}\mid s,a)\cdot\pi^{n}(a^{\prime}\mid s).$
To proceed with the proof, we fix $\bm{\lambda}^{n}$. For notational convenience, we omit $\bm{\lambda}^{n}$ from the argument of the policy $\pi^{n}$. We first fix $\pi^{n}$ and consider two mean field distributions $\mathcal{L}^{n},\mathcal{L}^{n^{\prime}}\in\mathcal{P}(\mathcal{S}\times\mathcal{A})$. Then we have
$\left|\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n},\pi^{n})(s^{\prime},a^{\prime})-\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n^{\prime}},\pi^{n})(s^{\prime},a^{\prime})\right|$ $=\left|\sum_{s,a}\left(\mathcal{L}^{n}(s,a)-\mathcal{L}^{n^{\prime}}(s,a)\right)P^{n}(s^{\prime}|s,a^{\prime})\pi^{n}(a^{\prime}|s)\right|$
$\leq\sum_{s,a}\left|\mathcal{L}^{n}(s,a)-\mathcal{L}^{n^{\prime}}(s,a)\right|P^{n}(s^{\prime}|s,a^{\prime})\pi^{n}(a^{\prime}|s).$
Now summing over all $(s^{\prime},a^{\prime})\in\mathcal{S}\times\mathcal{A}$ and using the fact that both $\pi^{n}(\cdot|s)$ and $P^{n}(\cdot|s,a^{\prime})$ are probability mass functions, we have that
$\|\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n},\pi^{n})-\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n^{\prime}},\pi^{n})\|_{1}=\sum_{s^{\prime},a^{\prime}}\left|\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n},\pi^{n})(s^{\prime},a^{\prime})-\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n^{\prime}},\pi^{n})(s^{\prime},a^{\prime})\right|$
$\leq\sum_{s^{\prime},a^{\prime}}\sum_{s,a}\left|\mathcal{L}^{n}(s,a)-\mathcal{L}^{n^{\prime}}(s,a)\right|P^{n}(s^{\prime}|s,a^{\prime})\pi^{n}(a^{\prime}|s)$
$=\sum_{s,a}\left|\mathcal{L}^{n}(s,a)-\mathcal{L}^{n^{\prime}}(s,a)\right|\sum_{a^{\prime}}\pi^{n}(a^{\prime}|s)\sum_{s^{\prime}}P^{n}(s^{\prime}|s,a^{\prime})=\sum_{s,a}\left|\mathcal{L}^{n}(s,a)-\mathcal{L}^{n^{\prime}}(s,a)\right|=\|\mathcal{L}^{n}-\mathcal{L}^{n^{\prime}}\|_{1}.$
Next, we fix $\mathcal{L}^{n}$, and let $\pi^{n},\pi^{n^{\prime}}$ be two policies. Then:
$\left|\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n},\pi^{n})(s^{\prime},a^{\prime})-\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n},\pi^{n^{\prime}})(s^{\prime},a^{\prime})\right|$ $=\left|\sum_{s,a}\mathcal{L}^{n}(s,a)P^{n}(s^{\prime}|s,a^{\prime})\left[\pi^{n}(a^{\prime}|s)-\pi^{n^{\prime}}(a^{\prime}|s)\right]\right|$
$=\sum_{s,a}\mathcal{L}^{n}(s,a)P^{n}(s^{\prime}|s,a^{\prime})\left|\pi^{n}(a^{\prime}|s)-\pi^{n^{\prime}}(a^{\prime}|s)\right|.$
Summing over all $(s^{\prime},a^{\prime})$ yields
$\|\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n},\pi^{n})-\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n},\pi^{n^{\prime}})\|_{1}=\sum_{s^{\prime},a^{\prime}}\sum_{s,a}\mathcal{L}^{n}(s,a)P^{n}(s^{\prime}|s,a^{\prime})\left|\pi^{n}(a^{\prime}|s)-\pi^{n^{\prime}}(a^{\prime}|s)\right|$
$=\sum_{s,a}\mathcal{L}^{n}(s,a)\sum_{a^{\prime}}\left|\pi^{n}(a^{\prime}|s)-\pi^{n^{\prime}}(a^{\prime}|s)\right|\sum_{s^{\prime}}P^{n}(s^{\prime}|s,a^{\prime})=\sum_{s,a}\mathcal{L}^{n}(s,a)\cdot\|\pi^{n}(\cdot|s)-\pi^{n^{\prime}}(\cdot|s)\|_{1}$
$\leq\sup_{s}\|\pi^{n}(\cdot|s)-\pi^{n^{\prime}}(\cdot|s)\|_{1}\cdot\sum_{s,a}\mathcal{L}^{n}(s,a)=\sup_{s}\|\pi^{n}(\cdot|s)-\pi^{n^{\prime}}(\cdot|s)\|_{1},$
where the last equality holds as the mean-field $\mathcal{L}^{n}(s,a)$, as defined in (9), is a joint probability distribution over $\mathcal{S}\times\mathcal{A}$. Combining the two results above, we have that
$\|\Gamma^{n}_{2}(\mathcal{L}^{n},\pi^{n})-\Gamma^{n}_{2}(\mathcal{L}^{n^{\prime}},\pi^{n^{\prime}})\|_{1}=(1-\zeta)\|\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n},\pi^{n})-\tilde{\Gamma}^{n}_{2}(\mathcal{L}^{n^{\prime}},\pi^{n^{\prime}})\|_{1}$
$\hskip 10.0pt\leq(1-\zeta)\left(\|\mathcal{L}^{n}-\mathcal{L}^{n^{\prime}}\|_{1}+\sup_{s}\|\pi^{n}(\cdot|s)-\pi^{n^{\prime}}(\cdot|s)\|_{1}\right),$
and we let $L_{2}=L_{3}=1-\zeta$. ∎
A.4   Proof of Theorem 3
Before presenting the theorem, we first recall the definition of a contraction mapping and the classical Banach fixed-point theorem:
Definition 4 (Contraction Mapping).
Let $(\mathcal{X},d)$ be a non-empty complete metric space. A map $T:\mathcal{X}\mapsto\mathcal{X}$ is a contraction mapping on $\mathcal{X}$ if $\;\forall x,y\in\mathcal{X}$, there exists $c\in[0,1)$ such that $d(T(x),T(y))\leq cd(x,y)$.
Then, we present the Banach fixed point theorem:
Theorem 4 (Banach Fixed-Point Theorem [25]).
Let $(\mathcal{X},d)$ be a non-empty complete metric space with a contraction mapping $T:\mathcal{X}\mapsto\mathcal{X}$. Then $T$ admits a unique fixed point $x^{*}$ in $\mathcal{X}$. That is, $T(x^{*})=x^{*}$.
[…]
Proposition 4 (Continuity of LMP).
[…]
Proof.
[…]
$|D^{n}_{t}-D^{n}_{t^{\prime}}|\leq\overline{E}_{n}\sum_{a,s}|a||\mathcal{L}^{n}_{t}(s,a)-\mathcal{L}^{n}_{t^{\prime}}(s,a)|\leq\overline{E}_{n}\|\mathcal{L}^{n}_{t}-\mathcal{L}^{n}_{t^{\prime}}\|_{1},$
[…]
Proof of Theorem 3. The key here is to show that the consistency operator $\Gamma^{n}_{2}$ is a contraction mapping. To do so, we focus on comparing the distributions at the same time of day across consecutive days. Let $H$ denote the length of a full day, and fix any time index $t\geq 0$. Due to the presence of diurnal patterns, it is natural to analyze the evolution of the mean-field distributions at times $t$ and $t+H$, corresponding to the same time period on two successive days. Let $\mathcal{L}^{n}_{t},\mathcal{L}^{n}_{t+H}\in\mathcal{P}(\mathcal{S}\times\mathcal{A})$ denote the mean-field distributions for aggregator $n$ at these two time points. We aim to establish that the update operator is contractive with respect to the $\ell_{1}$-norm over the joint state–action distribution:
$\|\Gamma^{n}_{2}(\mathcal{L}^{n}_{t},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t}))-\Gamma^{n}_{2}(\mathcal{L}^{n}_{t+H},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t+H}))\|_{1}{\leq}\|\Gamma^{n}_{2}(\mathcal{L}^{n}_{t},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t}))-\Gamma^{n}_{2}(\mathcal{L}^{n}_{t},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t+H}))\|_{1}$
$\hskip 200.0pt+\|\Gamma^{n}_{2}(\mathcal{L}^{n}_{t},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t+H}))-\Gamma^{n}_{2}(\mathcal{L}^{n}_{t+H},\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t+H}))\|_{1}$
$\hskip 5.0pt\overset{(a)}{\leq}L_{3}\|\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t})-\Gamma^{n}_{1}(s^{n},\bm{\lambda}^{n}_{t+H})\|_{1}+L_{2}\|\mathcal{L}^{n}_{t}-\mathcal{L}^{n}_{t+H}\|_{1}\overset{(b)}{\leq}L_{1}L_{3}\|\bm{\lambda}^{n}_{t}-\bm{\lambda}^{n}_{t+H}\|_{1}+L_{2}\|\mathcal{L}^{n}_{t}-\mathcal{L}^{n}_{t+H}\|_{1}$
=$\hskip 5.0pt\overset{(c)}{=}L_{1}L_{3}|\lambda^{n}_{t}-\lambda^{n}_{t+H}|+L_{2}\|\mathcal{L}^{n}_{t}-\mathcal{L}^{n}_{t+H}\|_{1}\overset{(d)}{\leq}(L_{1}L_{\text{MF}}L_{3}+L_{2})\|\mathcal{L}^{n}_{t}-\mathcal{L}^{n}_{t+H}\|_{1}.$
In the above derivations, inequality (a) follows from Theorem 2; while inequality (b) follows from Theorem 1. For the equality (c), it uses the fact that the only difference between the two vectors occurs at the $T_{\text{hour}}(t)$-th entry. The last inequality follows from the Lipschitz continuity of the LMPs as stated in Proposition 1. If the constant $L_{1}L_{\text{MF}}L_{3}+L_{2}\in[0,1)$, then the composition of the update operators defines a contraction mapping under the $\ell_{1}$-norm. Since $\lambda^{n}_{t}(\cdot)$ and the optimality operator $\Gamma^{n}_{1}$ are both single-valued, Banach’s fixed-point theorem guarantees the existence and uniqueness of the mean field equilibrium. $\Box$